# Multivariate Causal Effects: a Bayesian Causal Regression Factor Model

### v1

@@ Abstract.
The impact of wildfire smoke on air quality is a growing concern, contributing to air pollution through a complex mixture of chemical species with important implications for public health. While previous studies have primarily focused on its association with total particulate matter (pm2.5), the causal relationship between wildfire smoke and the chemical composition of pm2.5 remains largely unexplored. Exposure to these chemical mixtures plays a critical role in shaping public health, yet capturing their relationships requires advanced statistical methods capable of modeling the complex dependencies among chemical species. To fill this gap, we propose a Bayesian causal regression factor model that estimates the multivariate causal effects of wildfire smoke on the concentration of 27 chemical species in pm2.5 across the United States. Our approach introduces two key innovations: (i) a causal inference framework for multivariate potential outcomes, and (ii) a novel Bayesian factor model that employs a probit stick-breaking process as prior for treatment-specific factor scores. By focusing on factor scores, our method addresses the missing data challenge common in causal inference and enables a flexible, data-driven characterization of the latent factor structure, which is crucial to capture the complex correlation among multivariate outcomes. Through Monte Carlo simulations, we show the model’s accuracy in estimating the causal effects in multivariate outcomes and characterizing the treatment-specific latent structure. Finally, we apply our method to US air quality data, estimating the causal effect of wildfire smoke on 27 chemical species in pm2.5, providing a deeper understanding of their interdependencies.
@@ 1. Introduction
Wildfires have emerged as a critical environmental and public health concern, significantly deteriorating air quality through the emission of smoke and associated pollutants (Childs et al., 2022; Krasovich Southworth et al., 2025). In the United States, wildfire smoke exposure has increased significantly, with the burned area quadrupling over the past four decades—a trend expected to continue as climate conditions become increasingly favorable to fire activity. Wildfire smoke can rise high into the atmosphere and travel vast distances, affecting air pollution levels and health risks far from the source (Burke et al., 2021; O’Dell et al., 2021). As a results, it has received considerable attention in environmental health research (Reisen et al., 2015; Burke et al., 2021; Krasovich Southworth et al., 2025; Aguilera et al., 2021).
@@
While previous studies have focused on the impact of wildfire smoke on individual chemical components—such as phosphorus and nitrogen (Spencer and Hauer, 1991), lead (Odigie and Flegal, 2014), bromine and other metals (Li et al., 2023)—the recent work of Krasovich Southworth et al. (2025) has provided a more comprehensive analysis by quantifying the contribution of wildfire smoke to the concentration of 27 different chemical species in pm2.5 using a standard linear regression model. Their analysis primarily focuses on accounting for temporal variation and estimating the contribution of wildfire smoke to each chemical species in pm2.5. However, it does not jointly model the chemical species, limiting information sharing across species, and it is not adopting a causal inference prospective, restricting its ability to disentangle the causal effect of wildfire smoke.
@@
A causal inference framework is crucially needed to rigorously assess and quantify the effect of wildfire smoke on changes in chemical concentration levels, providing a deeper understanding of the associations highlighted in previous studies. Moreover, incorporating the high correlation structure among the different chemical species is critical to capturing deeper insights of the smoke composition and its broader environmental effects.
@@
To address these limitations, we propose a novel Bayesian factor model for causal inference, which leverages infinite mixture distributions as priors for the treatment-specific factor scores. The key novelty of our project lies in two main aspects. First, we define a causal inference framework in a Bayesian factor analysis model to account for the high correlation among multivariate potential outcomes, such as the chemical species in pm2.5. Second, we introduce a novel prior on the factor scores using a Dirichlet process (Quintana et al., 2022) with a probit linear model to capture treatment-specific heterogeneity in a data-driven way.
@@
Traditional Bayesian factor models have focused primarily on priors for factor loadings (Bhattacharya and Dunson, 2011; Schiavon et al., 2022; Frühwirth-Schnatter et al., 2024; De Vito et al., 2021), typically assuming standard normal distributions for factor scores. Recent works (Zorzetto et al., 2024; Bortolato and Canale, 2024) have begun exploring priors on factor scores, since they define the quantity that each subject has on the coresponding factor, making them essential in many applications. However, the causal inference framework in Bayesian factor models introduces new challenges due to the fundamental missing data problem of causal inference (Rubin, 1986). Our proposed prior addresses this by allowing covariate-dependent structure in treatment-specific scores, enabling better estimation of latent confounding patterns.
@@
The paper is organized as follows. Section 2 introduces the causal inference setup, including the estimands of interest, causal assumptions, and identification strategy. Section 3 introduces our Bayesian factor model for causal inference, highlighting the innovation in our prior specification for factor scores and its critical role in this context. In Section 4, we assess the model’s performance through simulations, comparing it to state-of-art approaches in causal inference and evaluating the importance of our proposed prior for the factor score compared to the factor model with the standard prior. In Section 5, we apply our approach to estimate the causal effect of the presence of wildfire smoke on the concentration of 27 chemical species in pm2.5 across the United States, using data from common sources of Krasovich Southworth et al. (2025): F.E.D. IMPROVE (2024) and Childs et al. (2022). We conclude the paper with a discussion and future directions in Section 6.
@@ 2. Causal setup
Let $i$ denote the study unit, with $i=1,\dots,n$, and $T_{i}\in\{0,1\}$ the binary treatment random variable with observed value $t_{i}$. Let also ${\bf X}_{i}=(X_{1i},\dots,X_{pi})^{T}$ denote a vector of $p$ observed covariates (confounders), which may be either continuous or discrete, and $\mathbf{Y}_{i}\in\mathbb{R}^{q}$ the multivariate outcome of dimension $q$, where the outcome components are potentially highly correlated, a structure exploited by our Bayesian causal factor model. Throughout the paper, bold letters denote vectors.
Following the definition of the Rubin Causal Model (Rubin, 1974), we invoke the Stable Unit Treatment Value Assumption (SUTVA; Rubin, 1980), which is a combination of two assumptions: no interference among the units, i.e. the potential outcome values from unit $i$ do not depend on the treatment applied to other units, and consistency, i.e. each observed outcome is the individual’s potential outcome under their treatment. Under SUTVA, we define the potential outcomes, for unit $i$, as $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$, where ${\bf Y}_{i}(t)=(Y_{i1}(t),\dots,Y_{iq}(t))^{T}$ for $t\in\{0,1\}$. Specifically, ${\bf Y}_{i}(0)\in{\mathds{R}}^{q}$ is the $q$-variate outcome when the unit $i$ is assigned to the control group and ${\bf Y}_{i}(1)\in{\mathds{R}}^{q}$ is the $q$-variate outcome when the unit $i$ is assigned to the treatment group. Under consistency, therefore the observed outcome can be written as:
@@
Moreover, we assume an unmeasured variable ${U}_{i}$ for each unit $i$, which affects the treatment $T_{i}$ and the potential outcome $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$, the last exclusively through latent factors $\{{\bf L}_{i0},{\bf L}_{i1}\}$. In particular, Figure 1 illustrates the three cases considered in our causal setup.
Several key clarifications are necessary, with a more detailed explanation provided in the following sections. First, although we refer to $U$ as a single random variable, it can be generalized to a function of multiple unmeasured variables. Second, the latent factors serve as ‘proxy’ variable that adjust the unmeasured variable, leveraging the correlation among the multivariate outcomes. Third, any relations between the unmeasured variable ${U}_{i}$ and the confounders ${\bf X}_{i}$ are considered, as reported in Figure 1.
Therefore, when $U$ affects the outcome only via the latent factors—as assumed in this paper—then this unmeasured information can be recovered through the latent structure estimated in our proposed Bayesian causal factor model, formally introduced in Section 3.
Figure 1. Graphical representation of the causal pathway assumed in this paper with treatment $T$, measured confounders $X$, unmeasured confounders $U$, latent factors $L$, and multivariate outcome $Y$. The three scenarios consider different relationships of unmeasured confounders $U$: (1) confounding only through an effect on latent factors $L$, (2) confounding as an effect of measured confounders $X$, (3) cause of measured confounders $X$. The simulation study investigates each of these scenarios.
@@
Our goal is to estimate the causal effect of the treatment on each element of the multivariate outcome. Therefore, we define the quantity of interest as the Sample Average Treatment Effects (SATE)
@@
where the average is taken across the units $\{1,\dots,n\}$ and SATE $\in\mathbb{R}^{q}$ is the vector of causal effects for the components of the outcome.
As common in causal inference, these causal estimands cannot be estimated directly from observed data due to the missing potential outcomes. To address this, we adopt the following assumptions and leverage the features of our proposed factor model, specifically the relationship of treatment-specific factor scores—i.e., ${\bf l}_{0}$ and ${\bf l}_{1}$—with measured and unmeasured variables. The first two assumptions are taken from standard causal inference literature, and the third is newly defined in the context of the assumed latent factor structure.
Positivity assumption. Each unit $i\in\{1,\dots,n\}$ has a non-zero probability o of receiving each treatment level,
@@
Conditional ignorability assumption. The assignment of treatment is random in each group of units characterized by observed covariates ${\bf x}$ and latent factors scores ${\bf l}_{0}$ and ${\bf l}_{1}$, such that for each unit $i\in\{1,\dots,n\}$
@@
Indirect unmeasured confounding assumption. The unmeasured variable $U$ affects the potential outcomes only indirectly, through the latent factor and the observed confounders, such as
@@
This assumption can be visualize in Figure 1 from the absence of a direct arrow between the $U$ and ${\bf Y}$.
@@
Under the three assumptions from previous Subsection, the causal estimands (i.e., functions of potential outcomes) can be re-expressed as statistical estimands (i.e., functions of random variables) and can then be estimated using the observed data.
@@
Invoking the previous assumptions, we can obtain the statistical estimand for SATEs as the following
@@
where the inner expectation ${\mathds{E}}[{\bf Y}\mid T=t,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]$, for each treatment level $t\in\{0,1\}$, will be estimated using our proposed factor model (2) and the joint probability of the latent factor score $\mathds{P}({\bf L}_{0},{\bf L}_{1}\mid{\bf X})$ modeled via prior distribution specified in (3).
A similar expression applies for the SATEk; see Appendix B for further details and full proof of Property 1.
Property 1 underlines the pivotal role of the probability distribution of treatment-specific factor scores, and thus the necessity of defining a flexible prior, such as the one we propose in in (3). Moreover, the inner expectation ${\mathds{E}}[{\bf Y}\mid T=t,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]$ can be rewritten as conditional on the unmeasured variable $U$, as follows:
@@
where the indirect unmeasured confounding assumption assumption is invoked for the second step. The involvement of the unmeasured variable $U$ in the conditional probability of factor scores $\mathds{P}(l_{0},l_{1}\mid u,x)$ leads to a careful choice of the distribution of this probability. We postpone the discussion to Section 3.
@@ 3. Bayesian Causal Factor Model
Due to the multivariate nature of the outcome, we propose a factor analysis based approach (Thurstone, 1931; Anderson-Cook, 2004) to model the potential outcomes given the treatment and the confounders. Rather than performing multiple univariate analyses, our factor analysis approach (i) improves statistical power by leveraging the correlation structure among the multivariate outcome, and (ii) facilitates the recovery of information from the unmeasured variable ${U}$.
Specifically, for each unit $i\in\{1,\dots,n\}$, the outcome vector ${\bf Y}_{i}$ follows a multivariate normal distribution such that
@@
where $\boldsymbol{\beta}_{tj}$ denotes the $j$-th row of the matrix ${\bf B}_{t}$ in (2), such that $\boldsymbol{\beta}_{tj}=\{\beta_{tj1},\dots,\beta_{tjp}\}$ and $\beta_{tjk}$ is the element $({\bf B}_{t})_{j,k}$ with $k\in\{1,\dots,p\}$. We consider independent priors such that $\Sigma_{m_{t}}$ and $\Sigma_{\beta_{t}}$ are diagonal matrices for each $t$. Alternative priors—e.g., mixtures of Gaussian distributions or spike and local prior—could be employed.
The Role of the DDP in the Factor Score
@@
In the literature, Bayesian statistical inference in factor analysis primarily focuses on prior for factor loading matrix, whereas the prior for the factor score matrix is typically non-informative and chosen as a standard Gaussian distribution, assuming independence across elements.
@@
However, this prior is unsuitable for handling missing data, as required in the potential outcomes framework for causal inference. Since accurate imputation of missing data is crucial, an informed choice of prior distribution is necessary. To address this, we propose a prior for the treatment-specific factor scores that takes advantage of the fundamental properties of the dependent Dirichlet process (DDP; Quintana et al., 2022). In particular, the DDPs are particular infinite mixture models—see details in the following Subsection—where our specific choice is to incorporate covariates into the weights to enhance the precision of imputation.
@@
This capability is crucial, as it allows for the modeling of complex probability distributions while leveraging covariates to fully characterize them. This distinct advantage has made DDPs a widely adopted and powerful tool in causal inference settings involving scalar outcomes. For example, Zorzetto et al. (2024) introduces the confounder-dependent Bayesian mixture model where the DPSB identifies groups in the population that characterize the heterogeneity in causal effects, the enriched Dirichlet process is used by Roy et al. (2018) and Roy et al. (2024), respectively, to use DDPs for missing in random covariates and mediation analysis, and Schwartz et al. (2011) and Zorzetto et al. (2024) specify a DDP prior for principal stratification.
In our factor model, we want to specify a DDP prior that allows us to capture heterogeneity at both the factor and treatment levels. The observed data inform the heterogeneity in the probability distribution of factor scores, facilitating the identification and estimation of data-driven clusters. Specifically, covariates determine the clustering of units through the mixture weights (see Eq. (5)), ensuring that units with similar covariate profiles share cluster-specific characteristics relevant for imputing missing potential outcomes.
@@
This data imputation process can be viewed as analogous to the matching process typically used in study design, where the criteria are driven by the covariates. However, our prior specification (Eq. (3)) does not replace the study design in observational studies, as further discussed in the application section.
@@
To align with the standard factor analysis literature, we define the continuous density function ${\mathcal{H}}(\cdot;\omega)$, with parameter space $\Omega$, as a Gaussian kernel, where $\omega=(\eta,1/\tau)$ are the position and scale parameters, respectively.
@@
for $t=\{0,1\}$ and $j=\{1,\dots,J_{t}\}$, where $\{\pi_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ and $\{\omega^{(tj)}_{r}\}_{l\geq 1}$ represent infinite sequences of random weights and random kernels’ parameters, respectively. Notably, both random sequences depend on the level of treatment $t$ and the factor $j$, while the weights also depend on the values of the confounders ${\bf x}_{i}$.
The sequence of random parameters $\{\omega_{r}^{(tj)}\}_{r\geq 1}$ is independent and identically distributed, where each $r$-th element of the sequence is the joint parameters of the $r$-th Gaussian distribution in the mixture model in (3), such that $\omega_{r}^{(tj)}=(\eta_{r}^{(tj)},1/\tau_{r}^{(tj)})$. Following the stick-breaking representation (Sethuraman, 1994), the sequence of dependent weights is defined as
@@
Among the dependent nonparametric processes, the Dependent Probit Stick-Breaking (DPSB) (Rodriguez and Dunson, 2011) defines the process $\{V_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ as follows
@@
where $\Phi(\cdot)$ is the Probit function and $\{a_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ has Gaussian distributions with mean a linear combination of the $p$ covariates ${\bf x}_{i}$.
The nature of the prior distribution for the treatment-specific factors (3) allows us to introduce a latent categorical variable $S_{itj}$, for each factor $j\in\{1,\dots,J_{t}\}$ treatment $t\in\{0,1\}$ unit $i\in\{1,\dots.n\}$, describing clusters of units defined by heterogeneous responses to the treatment level $t$. Assuming $\Pr(S_{itj}=r)=\pi_{r}^{(tj)}({\bf x}_{i})$, then the prior distribution of the treatment-specific factors (3) can be rewritten, conditional on $S_{itj}$, as
@@
The proposed prior for treatment-specific factor scores offers several strengths, as outlined in the previous Subsection. Additionally, it serves as a powerful tool for inferring information about the relationships among the multiple outcomes, as governed by unobserved features $U$. Specifically, the flexibility of the DDP distribution enables it to capture this unmeasured information while effectively controlling bias.
@@
Indeed, using multivariate normal properties, we can express the marginal distribution of treatment-specific factor scores ${\bf l}_{itj}$ and $U_{i}$, for each unit $i$, treatment $t$, and factor $j\in\{1,\dots,J_{t}\}$, so we can rewrite the marginal distribution of treatment-specific factor scores ${\bf l}_{itj}$ and $U_{i}$, for each unit $i$, treatment $t$, and factor $j\in\{1,\dots,J_{t}\}$, as follows:

### v2

@@ Abstract.
The impact of wildfire smoke on air quality is a growing concern, contributing to air pollution through a complex mixture of chemical species with important implications for public health. Although previous studies have focused mainly on its association with total particulate matter (pm2.5), the causal relationship between wildfire smoke and the chemical composition of pm2.5 remains largely unexplored. Exposure to these chemical mixtures plays a critical role in shaping public health, but capturing their relationships requires advanced statistical methods capable of modeling the complex dependencies among chemical species. To fill this gap, we propose a Bayesian causal regression factor model that estimates the multivariate causal effects of wildfire smoke on the concentration of 27 chemical species in pm2.5 across the United States. Our approach introduces two key innovations: (i) a causal inference framework for multivariate potential outcomes, and (ii) a novel Bayesian factor model that employs a probit stick-breaking process as prior for treatment-specific factor scores. By focusing on factor scores, our method addresses the missing data challenge common to causal inference and enables a flexible, data-driven characterization of the latent factor structure, which is crucial to capture the complex correlation between multivariate outcomes. Through Monte Carlo simulations, we show the accuracy of the model in estimating the causal effects in multivariate outcomes and characterizing the treatment-specific latent structure. Finally, we apply our method to US air quality data, estimating the causal effect of wildfire smoke on 27 chemical species in pm2.5, providing a deeper understanding of their interdependencies.
@@ 1. Introduction
Wildfires have emerged as a critical environmental and public health concern, significantly deteriorating air quality through the emission of smoke and associated pollutants (Childs et al., 2022; Krasovich Southworth et al., 2025). In the United States, wildfire smoke exposure has increased significantly, with the burned area quadrupling over the past four decades, a trend expected to continue as climate conditions become increasingly favorable to fire activity. Wildfire smoke can rise high into the atmosphere and travel long distances, affecting air pollution levels and health risks far from the source (Burke et al., 2021; O’Dell et al., 2021). As a result, wildfire pollution has received considerable attention in environmental health research (Reisen et al., 2015; Burke et al., 2021; Krasovich Southworth et al., 2025; Aguilera et al., 2021).
@@
While previous studies have focused on the impact of wildfire smoke on individual chemical components —such as phosphorus and nitrogen (Spencer and Hauer, 1991), lead (Odigie and Flegal, 2014), bromine and other metals (Li et al., 2023)— the recent work of Krasovich Southworth et al. (2025) has provided a more comprehensive analysis by quantifying the contribution of wildfire smoke to the concentration of 27 different chemical species in pm2.5 using a standard linear regression model. Their analysis focuses primarily on accounting for temporal variation and estimating the contribution of wildfire smoke to each chemical species in pm2.5. However, it does not jointly model chemical species, limiting information sharing across species, and does not adopt a causal inference perspective, restricting its ability to disentangle the causal effect of wildfire smoke.
@@
A causal inference framework is crucially needed to rigorously assess and quantify the effect of wildfire smoke on changes in chemical concentration levels, providing a deeper understanding of the associations highlighted in previous studies. Moreover, incorporating the high correlation structure among the different chemical species is critical to capturing deeper insights into the composition of smoke and its broader environmental effects.
@@
To address these limitations, we propose a novel Bayesian factor model for causal inference, which leverages infinite mixture distributions as priors for the treatment-specific factor scores. The key novelty of our project lies in two main aspects. First, we define a causal inference framework within a Bayesian factor analysis model, to account for the high correlation between multivariate potential outcomes, such as the chemical species in pm2.5. Second, we introduce a novel prior on the factor scores using a Dirichlet process (Quintana et al., 2022) with a probit linear model to capture treatment-specific heterogeneity in a data-driven way.
@@
Traditional Bayesian factor models have focused primarily on priors for factor loadings (Bhattacharya and Dunson, 2011; Schiavon et al., 2022; Frühwirth-Schnatter et al., 2024; De Vito et al., 2021), typically assuming standard normal distributions for factor scores. Recent works (Zorzetto et al., 2024; Bortolato and Canale, 2024) have begun exploring priors on factor scores, since they define the subject-specific contribution of each factor to the variation, making them essential in many applications. However, introducing a causal inference framework in Bayesian factor models poses new challenges due to the fundamental missing data problem of causal inference (Rubin, 1986). Our proposed prior addresses this by allowing a covariate-dependent structure in treatment-specific scores, enabling better estimation of latent confounding patterns. In particular, our prior leverages the flexibility of the Dependent Dirichlet Process (DDP) (Quintana et al., 2022), building on a growing body of work in causal inference that employs DDP priors (e.g., Roy et al., 2024; Hu et al., 2023; Zorzetto et al., 2024), though primarily in the context of scalar outcomes.
@@
The paper is organized as follows. Section 2 introduces the causal inference setup, including the estimands of interest, causal assumptions, and identification strategy. Section 3 introduces our Bayesian factor model for causal inference, highlighting the innovation in our prior specification for factor scores and its critical role in this context. In Section 4, we assess the performance of the model performance through simulations, comparing it to state-of-art approaches in causal inference and evaluating the importance of our proposed prior for the factor scores, compared to the factor model with the standard prior. In Section 5, we apply our approach to estimate the causal effect of the presence of wildfire smoke on the concentration of 27 chemical species in pm2.5 across the United States, using data from common sources of Krasovich Southworth et al. (2025): F.E.D. IMPROVE (2024) and Childs et al. (2022). We conclude the paper with a discussion and future directions in Section 6.
@@ 2. Causal setup
Notation and Overview
Let $i$ denote the study unit, with $i=1,\dots,n$. Foe each unit, let $T_{i}\in\{0,1\}$ the binary treatment random variable with observed value $t_{i}$, let ${\bf X}_{i}=(X_{1i},\dots,X_{pi})^{T}$ be the vector of $p$ observed covariates (confounders), which may be either continuous or discrete, and let $\mathbf{Y}_{i}\in\mathbb{R}^{q}$ be the multivariate outcome of dimension $q$. The outcome components are potentially highly correlated, a structure acknowledged by our Bayesian causal factor model. Throughout the paper, bold letters denote vectors.
Following the definition of the Rubin Causal Model (Rubin, 1974), we make the Stable Unit Treatment Value Assumption (SUTVA; Rubin, 1980), which is a combination of two assumptions: no interference among the units, i.e. the potential outcome values from unit $i$ do not depend on the treatment applied to other units, and consistency, i.e. each observed outcome is the individual’s potential outcome under their treatment. Under SUTVA, we define the potential outcomes, for unit $i$, as $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$, where ${\bf Y}_{i}(t)=(Y_{i1}(t),\dots,Y_{iq}(t))^{T}$ for $t\in\{0,1\}$. Specifically, ${\bf Y}_{i}(0)\in{\mathds{R}}^{q}$ is the $q$-variate outcome when the unit $i$ is assigned to the control group and ${\bf Y}_{i}(1)\in{\mathds{R}}^{q}$ is the $q$-variate outcome when the unit $i$ is assigned to the treatment group. Under consistency, therefore the observed outcome can be written as:
@@
To concisely define the correlated nature of the multidimensional outcome, for each unit $i$, we assume an unmeasured variable ${U}_{i}$ that affects the treatment $T_{i}$ and the potential outcome $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$. Under the assumption of our proposed Bayesian causal factor model, formally introduced in Section 3. Bayesian Causal Factor Model, we model both factual and counterfactual latent factors $\{{\bf L}_{i0},{\bf L}_{i1}\}$ that capture the indirect effect of $U_{i}$ on the potential outcomes $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$, as illustrated in Figure 1. Figure 1 shows the three versions of the resulting conditional independence structure considered in our model.
Figure 1. Graphical representation of the causal pathway assumed in this paper, with treatment $T$, measured confounders $X$, unmeasured confounder $U$, latent factors $L$, and multivariate outcome $Y$. The three scenarios consider different relationships for unmeasured confounder $U$: (1) confounding only through an effect on latent factors $L$, (2) confounding as an effect of measured confounders $X$ on $U$, (3) confounding because $U$ is a cause of measured confounders $X$. The simulation study investigates each of these scenarios.
@@
Our goal is to estimate the vector of causal effects of the treatment on each component of the multivariate outcome. Therefore, we define the quantities of interest as the $q$-dimensional Sample Average Treatment Effects (SATE)
@@
where the average is taken across the units $\{1,\dots,n\}$ and SATE $\in\mathbb{R}^{q}$.
As common in causal inference, these causal estimands cannot be estimated directly from observed data due to the missing potential outcomes. To address this, we adopt the following assumptions and leverage our proposed factor model, specifically the relationship of treatment-specific factor scores —i.e., ${\bf l}_{0}$ and ${\bf l}_{1}$—with measured and unmeasured variables. The first assumption is standard in the causal inference literature, the second is common in concept but adapted to our setting, while the third is newly defined for the context of the assumed latent factor structure.
Positivity. Each unit $i\in\{1,\dots,n\}$ has a non-zero probability of receiving each treatment level,
@@
Conditional ignorability. The assignment of treatment is random in each group of units characterized by observed covariates ${\bf x}$ and latent factors scores ${\bf l}_{0}$ and ${\bf l}_{1}$, such that for each unit $i\in\{1,\dots,n\}$
@@
Indirectness of unmeasured confounding effects. The unmeasured variable $U$ affects the potential outcomes only indirectly, through the latent factor and the observed confounders, such as
@@
This assumption can be visualized in Figure 1 as the absence of a direct arrow between the $U$ and ${\bf Y}$.
@@
Under the three assumptions of Section Causal Estimands and Identifying Assumptions, the causal estimands (i.e., functions of potential outcomes) can be expressed as statistical estimands (i.e., functions of random variables) and can then be estimated using observed data.
@@
If the three assumptions of Section Causal Estimands and Identifying Assumptions hold, the statistical estimand for SATEs is:
@@
where the inner expectation ${\mathds{E}}[{\bf Y}\mid T=t,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]$, for each treatment level $t\in\{0,1\}$, will be estimated using our proposed factor model (2) and the joint probability of the latent factor scores $\mathds{P}({\bf L}_{0},{\bf L}_{1}\mid{\bf X})$ modeled via prior distribution specified in (3).
Property 1 underlines the pivotal role of the probability distribution of treatment-specific factor scores and thus the necessity of defining a flexible prior, such as the one we propose in (3). Moreover, the inner expectation ${\mathds{E}}[{\bf Y}\mid T=t,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]$ can be rewritten with an additional integration over the unmeasured variable $U$, as follows:
@@
where the indirectness of unmeasured confounding assumption is used in the second step.
@@ 3. Bayesian Causal Factor Model
To address the multivariate outcome, we propose an approach that adopt a factor regression model (Carvalho et al., 2008; Avalos-Pacheco et al., 2022; De Vito and Avalos-Pacheco, 2023), which uses factor analysis to model the potential outcomes given the treatment and the confounders. Rather than performing multiple univariate analyses, our factor analysis approach (i) improves statistical power by leveraging the correlation structure in the outcome, and (ii) facilitates the recovery of information from the unmeasured variable ${U}$, under the three cases considered in Figure 1.
Specifically, we assume that for each unit $i\in\{1,\dots,n\}$, the outcome vector ${\bf Y}_{i}$, follows a treatment-specific probability density distribution $g_{t}(\cdot)$, with $t\in\{0,1\}$, where $g_{t}(\cdot)$ is modeled as a multivariate normal distribution
@@
where $\boldsymbol{\beta}_{tj}$ denotes the $j$-th row of the matrix ${\bf B}_{t}$ in (2), such that $\boldsymbol{\beta}_{tj}=\{\beta_{tj1},\dots,\beta_{tjp}\}$ and $\beta_{tjk}$ is the element $({\bf B}_{t})_{j,k}$ with $k\in\{1,\dots,p\}$. We assume that $\Sigma_{m_{t}}$ and $\Sigma_{\beta_{t}}$ are diagonal matrices for each $t$.
The Role of the DDP in the Factor Scores
@@
In the literature, Bayesian statistical inference in factor analysis primarily focuses on priors for the factor loading matrix, whereas the prior for the factor score matrix is typically chosen as a standard Gaussian distribution.
@@
However, this prior is not suitable for handling the imputation of the missing data, as required in the potential outcomes framework for causal inference. Since accurate imputation of missing data is crucial, an informed choice of prior distribution is necessary. To address this, in the following Section Depend Mixture Distribution for Treatment-Specific Factor Scores, we propose a prior distribution for the treatment-specific factor scores that takes advantage of the properties of the dependent Dirichlet process (DDP; Quintana et al., 2022).
@@
The flexibility of the DDP has made it widely adopted in causal inference settings involving scalar outcomes. For example, Zorzetto et al. (2024) adopt the probit stick-breaking process (Rodriguez and Dunson, 2011) to identifies subpopulations that characterize the heterogeneity in causal effects. Roy et al. (2018) and Roy et al. (2024) leverage the enriched Dirichlet process (Wade et al., 2011) estimate causal effects in the presence of missing-at-random covariates and for mediation analysis, respectively. In addition, Schwartz et al. (2011) and Zorzetto et al. (2024) specify a DDP prior within the principal stratification framework.
In our factor model, our DDP prior has two characteristics: (i) it can written as infinite mixture of Gaussian distributions, and (ii) the probability of assignment to each mixture component depends on the covariates $X$—called cluster allocation in the DDP literature. Moreover, we specify a DDP prior that allows us to capture heterogeneity at both the factor and treatment levels. In fact, treatment-specific factor scores quantify the score/importance that each unit expresses on the corresponding factor under the specific treatment level. In the environmental application discussed in Section 5. Environmental Application, each factor score of a specific chemical indicates its contribution to explain the correlation among chemicals within the corresponding latent factor (reflecting underlying sources of metal and organic emissions). Notably, each factor score can vary depending on the presence or absence of wildfire smoke, reflecting how wildfire exposure alters the chemical composition of the air.
The characteristic (i) defines a flexible distribution for each factor score that captures the heterogeneity in the factor scores through the identification of different clusters, while characteristic (ii) associates the cluster allocation to specific values of the covariates.
@@
This improves not only the estimation of factor scores compared with the standard Gaussian distribution as prior, but also their imputation under unobserved treatment levels. Specifically, for a given factor, the predictive posterior distribution of a factor score of a unit under the unobserved treatment level is similar to the posterior distributions of factor scores of units with similar covariate profiles, when observed under the same treatment. This data imputation process can be viewed as analogous to the matching process typically used in study design, where the criteria are driven by the covariates. However, our prior specification (Eq. (3)) does not replace the study design in observational studies, as further discussed in the application section.
@@
Consistently with standard factor analysis literature, we define the continuous density function ${\mathcal{H}}(\cdot;\omega)$, with parameter space $\Omega$, as a Gaussian kernel, where $\omega=(\eta,1/\tau)$ are the location and scale parameters, respectively.
@@
for $t=\{0,1\}$ and $j=\{1,\dots,J_{t}\}$, where $\{\pi_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ and $\{\omega^{(tj)}_{r}\}_{l\geq 1}$ represent infinite sequences of random weights and random kernels’ parameters, respectively. Notably, both random sequences depend on the level of treatment, $t$, and factor, $j$, while the weights also depend on the values ${\bf x}_{i}$ of the confounders.
The elements of the sequence of random parameters $\{\omega_{r}^{(tj)}\}_{r\geq 1}$ are independent and identically distributed, where the $r$-th element of the sequence is the joint parameters of the $r$-th Gaussian distribution in the mixture model in (3), so that $\omega_{r}^{(tj)}=(\eta_{r}^{(tj)},1/\tau_{r}^{(tj)})$. Following the stick-breaking representation (Sethuraman, 1994), the sequence of dependent weights is defined as
@@
Among the dependent nonparametric processes, the Dependent Probit Stick-Breaking (DPSB) (Rodriguez and Dunson, 2011) defines the process $\{V_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ as follows:
@@
where $\Phi(\cdot)$ is the Probit function and $\{a_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ have Gaussian distributions whose mean is a linear combination of the $p$ covariates ${\bf x}_{i}$.
This prior distribution for treatment-specific factors (3) allows us to introduce a latent categorical variable $S_{itj}$, for each factor $j\in\{1,\dots,J_{t}\}$ treatment $t\in\{0,1\}$ unit $i\in\{1,\dots.n\}$, describing clusters of units defined by heterogeneous responses to the treatment level $t$. Assuming $\Pr(S_{itj}=r)=\pi_{r}^{(tj)}({\bf x}_{i})$, then the prior distribution of the treatment-specific factors (3) can be rewritten, conditional on $S_{itj}$, as
@@
The proposed prior for treatment-specific factor scores offers several strengths, as outlined in Section The Role of the DDP in the Factor Scores. Additionally, it serves as a powerful tool for inferring information about the relationships among the multiple outcomes, as governed by unobserved features $U$. Specifically, the flexibility of the DDP distribution enables it to capture this unmeasured information while effectively controlling bias.
@@
Using multivariate normal properties, we can rewrite the marginal distribution of treatment-specific factor scores ${\bf l}_{itj}$ and $U_{i}$, for each unit $i$, treatment $t$, and factor $j\in\{1,\dots,J_{t}\}$, as follows: